# ProMQA: Question Answering Dataset for Multimodal Procedural Activity Understanding

### v1

@@ Abstract
Multimodal systems have great potential to assist humans in procedural activities, where people follow instructions to achieve their goals. Despite diverse application scenarios, systems are typically evaluated on traditional classification tasks, e.g., action recognition or temporal action segmentation. In this paper, we present a novel evaluation dataset, ProMQA, to measure system advancements in application-oriented scenarios. ProMQA consists of 401 multimodal procedural QA pairs on user recording of procedural activities coupled with their corresponding instruction. For QA annotation, we take a cost-effective human-LLM collaborative approach, where the existing annotation is augmented with LLM-generated QA pairs that are later verified by humans. We then provide the benchmark results to set the baseline performance on ProMQA. Our experiment reveals a significant gap between human performance and that of current systems, including competitive proprietary multimodal models. We hope our dataset sheds light on new aspects of models’ multimodal understanding capabilities.111Code and data are available at https://github.com/kimihiroh/promqa.
@@ 1 Introduction
Procedures are human knowledge of experience that enables one to obtain an expected outcome without much trial and error. Yet, following procedures (i.e., a set of instructions), itself requires skills such as, in cooking (Peddi et al., 2023), assembly (Sener et al., 2022), or surgery (Beyer-Berjot et al., 2016), among others. In supporting such user activities, current evolving multimodal foundation models like GPT-4o (OpenAI, 2024) and Claude 3.5 Sonnet (Anthropic, 2024) have great potential by monitoring the situation through the perception of a user’s wearable device. Despite such diverse application scenarios, existing studies typically provides traditional, but less practical evaluation testbeds. To support an application-oriented evaluation, we present a novel multimodal question-answering (QA) dataset for understanding procedural activity, annotated by our cost-effective human-LLM collaborative approach.
When supporting procedural activities, an assistant should comprehend information from multiple sources: 1) Actual process from their perception; 2) Each step and the overall flow from instructions. For instance, in cooking, answering “What is the next step now?” requires an assistant to recognize which steps have been completed until “now” from its video recording and identify what else/next should be done from its recipe. Assuming recipes are typically written in text, assistants receive multimodal information of how one did it as video and how one should do it as text. Prior work has explored the task in text-only, unimodal setting, where a user verbalizes all of their actions (Le et al., 2023). However, it is not ideal in practice as a beginning cook might give misleading explanations that cannot be corrected by a system without raw information (video) about the actual process.
@@
In this work, we present a novel dataset, ProMQA (Procedural Multimodal Question Answering), to evaluate models’ capabilities of understanding procedural activities in multimodal settings (§2). Our work is motivated by the fact that a well-adapted testbed is indispensable and can stimulate system development. In the dataset construction, we repurposed videos and recipes from the existing CaptainCook4D (Peddi et al., 2023) dataset. Then, for QA annotation, we employ a human-LLM collaborative approach, where LLMs first generate QA pairs and humans verify them to ensure the quality, inspired by the recent advances in synthetic data generation (Wang et al., 2023) (§3). While LLMs cost-effectively generate candidate QA pairs, the manual verification process ensures the quality of the resulting dataset. Specifically, among 500 generated QA pairs, around 80% were retained with additional human-written answers through the verification. Finally, to establish the baseline performance, we benchmark the following approaches: unimodal models, Socratic models (Zeng et al., 2022), and both open and proprietary multimodal models. Our benchmark experiments reveal that, while humans can reasonably perform the task, the dataset is challenging even for proprietary multimodal models, which show strong performance on other vision-language tasks (§4).
@@ 2 ProMQA
| Dataset Name | Multimodal | Video | Procedural | Explicit Instruction | QA | Open Vocab | LLM Scoring |
| YouCook2 (Zhou et al., 2018) | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ |
| How2QA (Li et al., 2020) | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ |
| Assembly101 (Sener et al., 2022) | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ |
| EgoSchema (Mangalam et al., 2023) | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
| CaptainCook4d (Peddi et al., 2023) | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ |
| GazeVQA (Ilaslan et al., 2023) | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
| MMBench (Liu et al., 2024b) | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ |
| OpenEQA (Majumdar et al., 2024) | ✓ | ✓ | ✗ | ✗ | ✓ | ✓ | ✓ |
@@
Our goal is to facilitate the development of procedural-activity support systems. ProMQA consists of 401 multimodal procedural QA pairs that require both recipes and video recordings to answer. It is constructed with our human-LLM collaborative approach on top of existing annotation (§3). In Table 1, we compare our dataset with similar multimodal datasets. Our dataset uniquely supports the assessment of multimodal procedural activity understanding as the QA task, which can serve as a complementary testbed to advance the model’s multimodal procedural activity understanding.
@@ 2.2 Multimodal Procedural QA
| Question type | Target | Example question |
| Process-level | ||
| Missing | Missing recipe steps | Did I miss any steps so far? |
| Next | Next recipe steps | What is the next step now? |
| Order | Errors w.r.t. recipe step ordering | Should I have done any steps in a different order? |
| Step-specific | ||
| Measurement | Errors in measurement (e.g., 2 cups instead of 1 cup) | Did I measure water correctly? |
| Preparation | Other errors in preparation (e.g., cilantro instead of oregano) | Did I add the correct spice? |
| Technique | Errors in cooking technique (e.g., chop instead of slice) | Did I prepare onion correctly? |
| Temperature | Errors in temperature (e.g., high instead of low) | Was the heat level correct? |
| Timing | Errors in duration (e.g., 2 min instead of 5 min) | Did I microwave it for long enough? |
@@
Among valid multimodal procedural questions, we categorize them into two groups, where each is further divided into specific question types. Process-level questions focus on multiple steps: missing, next, and order. Step-specific questions are questions about individual steps: measurement, preparation, technique, temperature, and timing. Examples of each type and their target phenomenon can be found in Table 3.
| #example (#question) | #distinct recipe | avg. #steps/ recipe | #distinct recording | avg. length of recording | avg. #steps/ recording | avg. #answers/ question | avg. #words/ question | avg. #words/ answer |
@@ 2.3 Statistics
We show the general statistics of our dataset in Table 4. In addition, among the 401 examples, 225 examples have no errors in previous steps (clean) and 176 examples have at least one error in previous steps. Figure 2 illustrates the high approved rate for questions, which obtained around 50% more additional human-written answers. We also compare the cost of our human-LLM collaborative approach and the estimate of the full-human annotation (Table 5). QA annotation for evaluation/test data typically consists of an initial QA creation, followed by verification to assure the quality. We only compare the cost of the QA creation/generation part as our annotation framework replaces humans with LLMs in the initial QA creation (§3). According to the table, our approach substantially reduces the cost of the QA creation part.
@@ 3 Annotation: Generate-then-Verify
In this work, we take a human-LLM collaborative approach to annotate QA pairs: LLMs generate QA pairs with lower cost, then humans verify them to ensure quality. We hypothesize that LLMs can substantially generate valid questions when given sufficient information, inspired by synthetic data generation (Wu et al., 2024). Specifically, we leverage existing annotations of action and error labels to form textual prompts. We note that, as our annotation framework is LLM agnostic, it can plug and play LLMs, and importantly, it can benefit from ongoing LLMs’ improvement.
@@ 3.1 Source & Preprocess
After obtaining approximately 11,000 examples from this process, we sampled 500 examples by taking the following points into account to increase diversity: (1) Sample at most one example for each question type from each recording, (2) Evenly sample examples with errors (noisy) and without errors (clean) in previous steps for all types, and (3) Evenly sample examples that do and do not have target recipe steps for next and missing types.333Example question without target recipe steps: “Did I miss any step?” “No.”
@@ 3.2 QA Generation
| Template | QA Generator | ||
| GPT-4o | Gemini 1.5 pro | Claude 3.5 Sonnet | |
| excerpt & step | 0.71 | 0.69 | 0.68 |
@@ 3.3 Verification
We first created an annotation guideline and hired 6 people with graduate degrees in NLP-related fields for our annotation. Among the participants, five people served as first-stage annotators, while the other, who was also involved in the guideline development, took the adjudicator’s role. This adjudicator performed the verification and human judgment in §3.2 and §4.2 as well. To help familiarize the annotators with the task, we conducted a training phase in which each annotator verified 20 examples and received personalized feedback. Following the training session, we initiated the main phase. On average, judgment agreements were for both questions and answers. After the verification, we obtained 401 verified examples .
@@ 4 Benchmarking
| Model | Avg. | w/ Error | Question Type | ||||||||
| clean | noisy | missing | next | order | measurement | preparation | technique | temperature | timing | ||
| Llama 3.1 70B Instruct | 31.5 | 33.2 | 30.2 | 35.5 | 38.3 | 12.5 | 2.9 | 14.3 | 5.6 | 30.0 | 20.0 |
| LLaVA 1.5 13B (50f, 288p) & Llama 3.1 70B Instruct | 37.5 | 38.9 | 36.4 | 41.9 | 45.3 | 12.5 | 11.8 | 14.3 | 11.1 | 50.0 | 18.0 |
| VideoLLaMA2 72B (8f, 336p) | 39.8 | 49.4 | 32.2 | 46.3 | 49.7 | 20.0 | 0.0 | 21.4 | 11.1 | 25.0 | 8.0 |
| Qwen2-VL 72B (100f, 336p) | 31.2 | 32.1 | 30.4 | 34.1 | 37.0 | 45.0 | 0.0 | 10.7 | 0.0 | 20.0 | 14.0 |
| GPT-4o (50f, 765p) | 40.4 | 39.5 | 41.1 | 39.9 | 45.9 | 45.0 | 29.4 | 17.9 | 27.8 | 55.0 | 24.0 |
| Gemini 1.5 Pro (50f, 765p) | 25.2 | 27.0 | 23.8 | 27.4 | 29.7 | 15.0 | 17.6 | 7.1 | 16.7 | 20.0 | 12.0 |
| Claude 3.5 Sonnet (10f, 765p) | 44.1 | 48.9 | 40.4 | 44.6 | 58.2 | 27.5 | 8.8 | 14.3 | 5.6 | 25.0 | 28.0 |
| Human | 74.5 | 83.5 | 65.8 | — | — | — | — | — | — | — | — |
| #choice | Context information | ||
| default | DOT | step | |
| binary | 0.67/0.86 | 0.57/0.80 | 0.71/0.89 |
| ternary | 0.67/0.69 | 0.58/0.64 | 0.76/0.75 |
| Template | Evaluator | ||
| GPT-4o | Claude 3.5 Sonnet | Gemini 1.5 Pro | |
| ternary & step | 0.83/0.82 | 0.79/0.77 | 0.66/0.68 |
@@ 4.2 LLM-as-a-Judge
We aim to identify a prompt template and an LLM that yields a high correlation with human judges. We consider two key aspects in templates: 1) the number of choices in the Likert scale and 2) the context information. For choices, we consider “binary” (match and unmatch) and “ternary” (match, partial-match, and unmatch). For context, we examine three settings: With a question, gold answers, and a predicted answer as the fundamental elements (“default”), we then incorporate either instruction (“DOT”) or recordings as step descriptions (“step”). Candidate evaluators include GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro. In the experiment, we feed inputs based on the verified examples in §3.2 to LLM-evaluators to obtain predictions. Then, we obtain judgments from these LLMs with all combinations. As a comparison, we obtain human judgments, where one person judges the predictions with both binary or ternary options. We consider Pearson correlation coefficient (Pearson, 1895) and match accuracy as our metrics. Table 9 shows the average scores across three evaluator models. We found that the combination of “ternary” and “step” produces the highest correlation. With the best combination, we compare the evaluators. Table 10 shows that GPT-4o has the best correlation with the human judgments. In the benchmark experiment, we scaled judgment scores from 0-2 to 0-100 by multiplying 50.

### v2

@@ Abstract
Multimodal systems have great potential to assist humans in procedural activities, where people follow instructions to achieve their goals. Despite diverse application scenarios, systems are typically evaluated on traditional classification tasks, e.g., action recognition or temporal action segmentation. In this paper, we present a novel evaluation dataset, ProMQA, to measure system advancements in application-oriented scenarios. ProMQA consists of 401 multimodal procedural QA pairs on user recording of procedural activities, i.e., cooking, coupled with their corresponding instructions/recipes. For QA annotation, we take a cost-effective human-LLM collaborative approach, where the existing annotation is augmented with LLM-generated QA pairs that are later verified by humans. We then provide the benchmark results to set the baseline performance on ProMQA. Our experiment reveals a significant gap between human performance and that of current systems, including competitive proprietary multimodal models. We hope our dataset sheds light on new aspects of models’ multimodal understanding capabilities.111Code and data are available at https://github.com/kimihiroh/promqa.
@@ 1 Introduction
Procedures are human knowledge of experience that enables one to obtain an expected outcome without much trial and error. Yet, following procedures (i.e., a set of instructions), itself requires skills such as, in cooking (Peddi et al., 2023), assembly (Sener et al., 2022), or surgery (Beyer-Berjot et al., 2016), among others. In supporting such user activities, current evolving multimodal foundation models like GPT-4o (OpenAI, 2024) and Claude 3.5 Sonnet (Anthropic, 2024) have great potential by monitoring the situation through the perception of a user’s wearable device. Despite such diverse application scenarios, existing studies typically provide traditional, but less practical evaluation testbeds. To support an application-oriented evaluation, we present a novel multimodal question-answering (QA) dataset for understanding procedural activity, produced by our cost-effective human-LLM collaborative approach.
When supporting procedural activities, an assistant should comprehend information from multiple sources: 1) Actual process from their perception; 2) Each step and the overall flow from instructions. For instance, in cooking, answering “What is the next step now?” requires an assistant to recognize which steps have been completed until “now” from its video recording and identify what else/next should be done from its recipe. Assuming recipes are typically written in text, assistants receive multimodal information of how one did it as video and how one should do it as text. Prior work has explored the task in a text-only, unimodal setting, where a user verbalizes all of their actions (Le et al., 2023). However, it is not ideal in practice as a beginning cook might give misleading explanations that cannot be corrected by a system without raw information (video) about the actual process.
@@
In this work, we present a novel dataset, ProMQA (Procedural Multimodal Question Answering), to evaluate models’ capabilities of understanding procedural activities in multimodal settings (§2). Our work is motivated by the fact that a well-adapted testbed is indispensable and can stimulate system development. In the dataset construction, we repurposed videos and recipes from the existing CaptainCook4D (Peddi et al., 2023) dataset. Then, for QA annotation, we employ a human-LLM collaborative approach, where LLMs first generate QA pairs and humans verify them to ensure the quality, inspired by the recent advances in synthetic data generation (Mangalam et al., 2023) (§3). While LLMs cost-effectively generate candidate QA pairs, the manual verification process ensures the quality of the resulting dataset. Specifically, among 500 generated QA pairs, around 80% were retained with additional human-written answers through the verification. Finally, to establish the baseline performance, we benchmark the following approaches: unimodal models, Socratic models (Zeng et al., 2022), and both open and proprietary multimodal models. Our benchmark experiments reveal that, while humans can reasonably perform the task, the dataset is challenging even for proprietary multimodal models that show strong performance on other vision-language tasks (§4).
@@ 2 ProMQA
Multimodal | Video | Procedural |
QA |
| Assembly101 (Sener et al., 2022) | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ||||||||
| IndustReal (Schoonbeek et al., 2024) | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ||||||||
| YouCook2 (Zhou et al., 2018) | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ||||||||
| CaptainCook4d (Peddi et al., 2023) | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ||||||||
| How2QA (Li et al., 2020) | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ||||||||
| MMBench (Liu et al., 2024b) | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ | ||||||||
| EgoSchema (Mangalam et al., 2023) | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ||||||||
| GazeVQA (Ilaslan et al., 2023) | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ||||||||
| OpenEQA (Majumdar et al., 2024) | ✓ | ✓ | ✗ | ✗ | ✓ | ✓ | ✓ | ||||||||
@@
| Question type | Target | Example question | |
| Process-level | |||
| Missing | Missing recipe steps | Did I miss any steps so far? | |
| Next | Next recipe steps | What is the next step now? | |
| Order | Errors w.r.t. recipe step ordering | Should I have done any steps in a different order? | |
| Step-specific | |||
| Measurement |
Did I measure water correctly? | |
| Preparation |
Did I add the correct spice? | |
| Technique |
Did I prepare onion correctly? | |
| Temperature |
Was the heat level correct? | |
| Timing |
Did I microwave it for long enough? | |
Our goal is to facilitate the development of procedural-activity support systems. ProMQA consists of 401 multimodal procedural QA pairs that require both recipes and video recordings to answer. It is constructed with our human-LLM collaborative approach on top of existing cooking recording and annotation (§3). In Table 1, we compare our dataset with similar multimodal datasets. Our dataset uniquely supports the assessment of multimodal procedural activity understanding as the QA task, which can serve as a testbed to advance the model’s multimodal procedural activity understanding.
@@ 2.2 Multimodal Procedural QA
Among valid multimodal procedural questions, we categorize them into two groups, where each is further divided into specific question types, following CaptainCook4D. Process-level questions focus on multiple steps: missing, next, and order. Step-specific questions are questions about individual steps: measurement, preparation, technique, temperature, and timing. Examples of each type and their descriptions can be found in Table 3.
@@ 2.3 Statistics
We show the general statistics of our dataset in Table 4. Among the 401 examples, 225 examples have no errors in previous steps (clean) and 176 examples have at least one error in previous steps. Figure 2 illustrates the high approval rate for questions, while approximately 50% of answers were added by humans through the verification process. In addition to showing the total count of each answer characteristic, we also count the number of examples with each combination of answer sources and types, as shown in Figure 3 and 4. For these analyses, while we used the answer source information retained through the annotation process, we obtained the answer type information by asking one annotator to categorize each answer into one of three types. We further compare the cost of our human-LLM collaborative approach and the estimate of the full-human annotation in Table 5. QA annotation for evaluation/test data typically consists of two steps: initial QA creation, followed by verification to assure the quality. We only compare the cost of the QA creation/generation part as our annotation framework replaces humans with LLMs in the initial QA creation (§3). According to the table, our approach substantially reduces the cost of the QA creation part.
@@ 3 Annotation: Generate-then-Verify
In this work, we take a human-LLM collaborative approach to annotate QA pairs: LLMs generate QA pairs with lower cost, then humans verify them to ensure quality. We hypothesize that LLMs can substantially generate valid questions when given sufficient information, inspired by synthetic data generation (Mangalam et al., 2023; Wu et al., 2024). Specifically, we leverage existing annotations of action and error labels to form textual prompts. We note that, as our annotation framework is LLM agnostic, it can plug and play LLMs, and importantly, it can benefit from ongoing LLMs’ improvement.
@@ 3.1 Source & Preprocess
After obtaining approximately 11,000 examples from this process, we sampled 500 examples by taking the following points into account to increase diversity: (1) Sample one example for each question type from each recording; (2) Evenly sample examples with errors (noisy) and without errors (clean) in previous steps for all types; (3) Evenly sample examples that do and do not have target recipe steps for next and missing types.333Example question without target recipe steps: “Did I miss any step?” “No.” Note that the activities in CaptainCook4D do not always result in the expected outcomes, i.e., failed procedures are included. Hence, our noisy examples allow us to generate QAs on top of unaddressed and/or irreparable errors from previous steps.
@@ 3.2 QA Generation
| Template | QA Generator | |||
| GPT-4o |
| excerpt & step | 0.71 | 0.69 | 0.68 | |
@@ 3.3 Verification
We first created an annotation guideline and hired 6 people with graduate degrees in NLP-related fields for our annotation. Among the participants, five people served as first-stage annotators, while the other, who was also involved in the guideline development, took the adjudicator’s role. This adjudicator performed the answer categorization, verification, and human judgment in §2.3, §3.2, and §4.2 as well. To help familiarize the annotators with the task, we conducted a training phase in which each annotator verified 20 examples and received personalized feedback. Following the training session, we initiated the main phase. On average, judgment agreements were for both questions and answers. After the verification, we obtained 401 verified examples .
@@ 4.1 Target Models
| #choice | Context information | ||
| default | DOT | step | |
| binary | 0.67/0.86 | 0.57/0.80 | 0.71/0.89 |
| ternary | 0.67/0.69 | 0.58/0.64 | 0.76/0.75 |
| Template | Evaluator | |||
| GPT-4o |
| ternary & step | 0.83/0.82 | 0.79/0.77 | 0.66/0.68 | |
| Model | Avg. | Error | Question Type | ||||||||||
| clean | noisy | missing | next | order | measurement | preparation | technique | temperature | timing | ||||
| Llama 3.1 70B Instruct | 31.5 | 33.2 | 30.2 | 35.5 | 38.3 | 12.5 | 2.9 | 14.3 | 5.6 | 30.0 | 20.0 | ||
37.5 | 38.9 | 36.4 | 41.9 | 45.3 | 12.5 | 11.8 | 14.3 | 11.1 | 50.0 | 18.0 | ||
| VideoLLaMA2 72B (8f, 336p) | 39.8 | 49.4 | 32.2 | 46.3 | 49.7 | 20.0 | 0.0 | 21.4 | 11.1 | 25.0 | 8.0 | ||
| Qwen2-VL 72B (100f, 336p) | 31.2 | 32.1 | 30.4 | 34.1 | 37.0 | 45.0 | 0.0 | 10.7 | 0.0 | 20.0 | 14.0 | ||
| GPT-4o (50f, 765p) | 40.4 | 39.5 | 41.1 | 39.9 | 45.9 | 45.0 | 29.4 | 17.9 | 27.8 | 55.0 | 24.0 | ||
| Gemini 1.5 Pro (50f, 765p) | 25.2 | 27.0 | 23.8 | 27.4 | 29.7 | 15.0 | 17.6 | 7.1 | 16.7 | 20.0 | 12.0 | ||
| Claude 3.5 Sonnet (10f, 765p) | 44.1 | 48.9 | 40.4 | 44.6 | 58.2 | 27.5 | 8.8 | 14.3 | 5.6 | 25.0 | 28.0 | ||
| Human* | (74.5) | (83.5) | (65.8) | — | — | — | — | — | — | — | — | ||
| Model | Avg. | Answer Source | Answer Type | ||||||||
| machine | human | both | direct | direct & suggestion | direct & intervention | suggestion | all | ||||
| Llama 3.1 70B Instruct | 31.5 | 33.4 | 28.0 | 30.5 | 31.7 | 30.8 | 34.1 | 25.0 | 0.0 | ||
37.5 | 40.7 | 32.9 | 34.8 | 38.8 | 34.6 | 34.1 | 12.5 | 0.0 | ||
| VideoLLaMA2 72B (8f, 336p) | 39.8 | 45.1 | 32.9 | 34.3 | 41.7 | 34.6 | 31.8 | 12.5 | 0.0 | ||
| Qwen2-VL 72B (100f, 336p) | 31.2 | 35.0 | 22.0 | 30.5 | 30.8 | 33.3 | 27.3 | 37.5 | 100.0 | ||
| GPT-4o (50f, 765p) | 40.4 | 41.4 | 29.3 | 47.1 | 40.2 | 39.7 | 52.3 | 18.7 | 50.0 | ||
| Gemini 1.5 Pro (50f, 765p) | 25.2 | 27.1 | 18.3 | 26.7 | 24.8 | 25.6 | 34.1 | 12.5 | 50.0 | ||
| Claude 3.5 Sonnet (10f, 765p) | 44.1 | 48.1 | 35.4 | 42.9 | 45.0 | 42.3 | 43.2 | 25.0 | 0.0 | ||
@@ 4.2 LLM-as-a-Judge
We aim to identify a prompt template and an LLM that yields a high correlation with human judges. We consider two key aspects in templates: 1) the number of choices in the Likert scale and 2) the context information. For choices, we consider “binary” (match and unmatch) and “ternary” (match, partial-match, and unmatch). For context, we examine three settings: With a question, gold answers, and a predicted answer as the fundamental elements (“default”), we then incorporate either instruction (“DOT”) or step descriptions from recordings (“step”). Candidate evaluators include GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro. In the experiment, we feed inputs based on the verified examples in §3.2 to LLM-evaluators to obtain predictions. Then, we obtain judgments from these LLMs with all combinations. As a comparison, we obtain human judgments, where one person judges the predictions with both binary or ternary options. We consider Pearson correlation coefficient (Pearson, 1895) and match accuracy as our metrics. Table 8 shows the average scores across three evaluator models. We found that the combination of “ternary” and “step” produces the highest correlation. With the best combination, we compare the evaluators. Table 9 shows that GPT-4o has the best correlation with the human judgments. In the benchmark experiment, we scaled judgment scores from 0-2 to 0-100 by multiplying 50.